OpenStudy (anonymous):

(2√5+3√7)^2 please show steps to solving this

sam (.sam.):

(2sqrt(5)+3sqrt(7))^(2) (2sqrt(5)+3sqrt(7))(2sqrt(5)+3sqrt(7)) (4sqrt(5)^(2)+12sqrt(5)sqrt(7)+9sqrt(7)^(2)) 4sqrt(5)^(2)+12sqrt(5)sqrt(7)+9sqrt(7)^(2) 20+12sqrt(5)sqrt(7)+9(7) 20+12sqrt(5)sqrt(7)+63 83+12sqrt(5)sqrt(7) 83+12sqrt35

OpenStudy (anonymous):

use this formula...\[\LARGE (a+b)^2=a^2+2ab+b^2\] we have... \[\LARGE (2\sqrt5+3\sqrt7)^2\] so... \[\Large (2\sqrt5+3\sqrt7)^2=(2\sqrt5)^2+2\cdot(2\sqrt5\cdot3\sqrt7)+(3\sqrt7)^2\] that becomes... \[\Large 4\cdot 5+2\cdot2\cdot3\cdot\sqrt{5\cdot7}+9\cdot7\] \[\LARGE 20+12\sqrt{35}+63\] \[\LARGE 83+12\sqrt{35}\]

OpenStudy (anonymous):

@HelloKalii if you don't understand it just let me know where you're stuck ;)